# ESPFormer: Doubly-Stochastic Attention with Expected Sliced Transport Plans

### v1

@@ 2.5 Soft Sorting
Calculating the sliced Wasserstein distance involves evaluating the quantile functions of the distributions, which, in the discrete case, can be boiled down to the sorting operation. Sorting is one of the most common operations in computer science. Yet, the piecewise-linear sorted value function and the integer-valued rank/argsort operators pose a significant obstacle for gradient-based optimization techniques, which are essential in deep learning, as neither of them is differentiable. To incorporate sorting operations into the backpropagation framework, differentiable approximations, known as soft sorting, have been explored. Examples include smoothed rank operators by adding Gaussian noise [47] and by using sigmoid surrogate functions [39], parameterizing permutations in terms of a differentiable relaxation [32], and relaxing the permutation matrices to be only row-wise stochastic [9]. Of note, Cuturi et al. [5] propose a differentiable proxy by viewing sorting as an optimal assignment problem and relaxing it to an optimal transport problem from the input values to an auxiliary probability measure supported on an increasing family of target values.
Prillo and Eisenschlos [38] introduce a simple yet highly effective continuous relaxation of the argsort operator using the softmax function: Given a vector $v\in\mathbb{R}^{N}$,
@@
where the softmax operator is applied row-wise, $d(\cdot,\cdot)$ can be any differentiable semi-metric, and $t$ is a temperature parameter that controls the degree of the approximation.
@@ 3.1 ESP for Uniform Discrete Distributions
Given our application of interest, we specifically focus on uniformly distributed discrete distributions with an equal number of support points. For a given $N\in\mathbb{N}$, denote the space of uniform discrete distributions with $N$ support as
@@
Consider the 1-dimensional slice of $\mu^{1}$ and $\mu^{2}$ in the $\theta$ direction: $\theta_{\#}\mu^{1}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\theta\cdot x_{i}}$ and $\theta_{\#}\mu^{2}=\frac{1}{N}\sum_{j=1}^{N}\delta_{\theta\cdot y_{j}}$, with $X\theta=[\theta\cdot x_{1},\theta\cdot x_{2},\cdots,\theta\cdot x_{N}]^{T}\in\mathbb{R}^{N}$ and $Y\theta=[\theta\cdot y_{1},\theta\cdot y_{2},\cdots,\theta\cdot y_{N}]^{T}\in\mathbb{R}^{N}$. There exists $A,B\in\mathbf{S}_{N}$ such that $AX\theta\in\mathbb{R}^{N}$ and $BY\theta\in\mathbb{R}^{N}$ are in sorted order, i.e.,
@@
or equivalently, $A^{T}B$ represents the transport map from $X\theta$ to $Y\theta$, i.e.,
@@
where $\tau$ denotes the ‘inverse temperature” hyperparameter, and $\sigma_{l}^{\tau}$ is defined as
@@
$\mathcal{D}^{p}_{p}(\mu^{1},\mu^{2};\theta)=\sum_{i=1}^{N}\sum_{j=1}^{N}\|x_{i}-y_{j}\|^{p}U_{\theta}(\{(x_{i},y_{j})\}).$ (10)
When $\tau=0$, $\sigma_{l}=\frac{1}{L}$ for all $l\in[1,2,\cdots,L]$, then $G$ is simply the mean of all the lifted plans.
@@ 3.2 ESP Attention
To bridge the abstract formulation of the Expected Sliced Plan (ESP) with its practical use in attention mechanisms, we interpret the query and key matrices as empirical probability measures due to their permutation-equivalent nature. Following our notation in Section 2.1, let $W_{Q},W_{K}\in\mathbb{R}^{m\times d},W_{V}\in\mathbb{R}^{d\times d}$ denote the query, key, and value matrices, respectively, and let $X=[x_{1},x_{2},\cdots,x_{N}],x_{i}\in\mathbb{R}^{d},\forall i$ denote the input tokens. Also, let $Q=W_{Q}X\in\mathbb{R}^{m\times N}$ and $K=W_{K}X\in\mathbb{R}^{m\times N}$ denote the queries and keys, respectively. Then, we think of doubly-stochastic attention as a transport plan between the empirical measures $\mu^{Q}$ and $\mu^{K}$, defined as
@@
A key consideration is the choice of slices, $\Theta$. While classic sliced OT samples slicers uniformly from $\mathbb{S}^{m-1}$, prior work suggests learning them [6, 33]. Though this approach is common, learning $\Theta$ introduces additional parameters, increasing the total count in ESP attention and complicating fair comparisons with other mechanisms. Notably, since the input distributions (keys and queries) are themselves learned, optimizing $\Theta$ may be unnecessary. Instead, the distributions can adapt to a fixed slicing scheme, as done in prior work like FSPool [59]. To avoid extra parameters, in this work, we propose using axis-aligned slices by setting $\Theta=\mathds{I}_{m\times m}$, the identity matrix. This leads to
@@
where $V=W_{V}X\in\mathbb{R}^{d\times N}$ is the value matrix. Figure 1 illustrates the construction of the ESP matrix $G$, which serves as the proposed transport-based attention map, directing how information flows from $V$ to the output. Moreover, Figure 2 compares classic attention (Sinkhorn $S^{0}$), Sinkhorn attention from [44] across different iteration counts, and our ESP attention under varying ‘inverse temperature” hyperparameters. As can be seen, while ESP provides a balanced distribution of attention due to its double stochasticity, the inverse temperature” parameter controls the number of other tokens each token should pay attention to. The overall pipeline of ESPFormer can be found in Algorithm 1.
@@
0:  Query matrix $Q\in\mathbb{R}^{m\times N}$, Key matrix $K\in\mathbb{R}^{m\times N}$,Value Matrix $V\in\mathbb{R}^{d\times N}$, SoftSort hyperparameter $t$,and ‘inverse temperature” hyperparameter $\tau$.
@@ 3.3 Runtime Complexity
ESPFormer begins with query and key projections ($Q=W_{Q}X$ and $K=W_{K}X$), each requiring $\mathcal{O}(mNd)$ operations, where $d$ corresponds to the output dimension. The subsequent SoftSort operation, which involves computing pairwise distances, incurs a complexity of $\mathcal{O}(N^{2})$ per slice. When applied across all $m$ slices, this yields a complexity of $\mathcal{O}(mN^{2})$. The computation of transport plans through matrix multiplications with soft permutation matrices contributes an additional $\mathcal{O}(mN^{2})$ operations. The final aggregation of transport plans requires summing $m$ plans of size $N\times N$, also contributing $\mathcal{O}(mN^{2})$ operations. Therefore, the overall runtime complexity of ESPFormer is $\mathcal{O}(mN(N+d))$. In comparison, Sinkhorn’s algorithm exhibits a runtime complexity of $\mathcal{O}((S+m)N^{2})$, where $S$ denotes the number of iterations. A key distinction lies in the parallelization capabilities: while Sinkhorn necessitates sequential processing over $S$ iterations, ESPFormer enables parallel computation across the $m$ slices. This parallelizability allows ESPFormer to scale efficiently with increasing $m$. As demonstrated in Figure 3, ESPFormer achieves superior runtime compared to Sinkformer. The wall clock runtime analysis of ESPFormer compared to all the baselines can be found in Appendix A.
Figure 3: Runtime analysis of ESPFormer versus Sinkformer with varying iteration counts ($S$) for sequence lengths $N\in\{50,100,500,1000\}$ with $d=1024$, averaged over 10 runs. ESPFormer demonstrates superior computational efficiency across all sequence lengths while preserving model expressivity.
@@ 4 Experiments
In this section, we evaluate the performance of ESPFormer across a diverse set of domains, including image classification using ViTs [7], point cloud classification using set transformers [24] and point cloud transformers [10], as well as sentiment analysis and neural machine translation via Transformers, highlighting the applicability of our approach in comparison to Differential Transformer [55], Sinkformer [44], and the Vanilla Transformer [51]. The details of our experimental setup and implementation can be found in Appendices B and C. [^1] Our implementation code will be released upon acceptance.
@@ 4.1 ModelNet 40 Classification
The ModelNet40 dataset [53] comprises 40 widely recognized 3D object categories and serves as a standard benchmark for point cloud classification. Transformers designed for point clouds and sets have been extensively evaluated on ModelNet40, with notable examples including Set Transformers [24] and Point Cloud Transformers [10]. Table 2 presents the classification results across four runs, comparing different attention mechanisms integrated into Set Transformers and Point Cloud Transformers. Notably, ESPFormer outperforms competing methods, demonstrating superior performance.
Table 1: Test accuracy for Sentiment Analysis on IMDb dataset.
@@
ESPFormer 85.50% 85.50% 85.47% 85.40%
Sinkformer 85.40% 85.40% 85.37% 85.3%
DiffTransformer 85.50% 85.45% 85.45% 85.40%
Transformer 85.30% 85.25% 85.25% 85.30%
Table 2: Test accuracy on the ModelNet40 dataset over four runs. Accuracies marked with are reported from [44].
@@
Set ESPFormer 89.6% 89.5% 89.4% 89.1%
Set Sinkformer 89.1% 88.4% 88.3% 88.1%
Set DiffTransformer 89.0% 88.7% 88.7% 88.6%
Set Transformer 87.8% 86.3% 85.8% 84.7%
Point Cloud ESPFormer 93.2% 92.9% 92.7% 92.6%
Point Cloud Sinkformer 93.1% 92.8% 92.7% 92.5%
Point Cloud DiffTransformer 93.1% 92.8% 92.7% 92.6%
Point Cloud Transformer 93.2% 92.5% 92.5% 92.3%
4.2 Sentiment Analysis
We next evaluate ESPFormer on the IMDB dataset [29] for sentiment analysis. Following Sander et al. [44], our architecture comprises an attention-based encoder followed by a max-pooling layer trained to classify movie reviews as either positive or negative. Table 1 presents the accuracy improvements achieved by using ESPFormer as the core attention module, compared to baseline models, showcasing its robustness in text classification.
@@ 4.3 Neural Machine Translation
We evaluate ESPFormer and Sinkformer using two reference models: the Transformer and its DiffTransformer counterpart [55]. Both models are trained using the fairseq sequence modeling toolkit [36] on the IWSLT’14 German-to-English dataset [2]. The architecture of both models consists of an encoder and a decoder, each with a depth of 6 layers. Initially, we trained both models for 25 epochs using the standard training procedure. After this phase, we performed a Plug-and-Play evaluation, where the attention heads were plugged into the pre-trained models and evaluated on their performance, as shown in Table 3. In this phase, we tested ESPFormer and Sinkformer attention heads, denoted by the respective models in the table, comparing their performance to the base Transformer and DiffTransformer models. Following the Plug-and-Play evaluation, we performed a Fine-Tune Boost phase, where the models were further fine-tuned for an additional 10 epochs. The fine-tuning led to further performance gains, as observed in the table, with ESPFormer showing the highest improvement in both the Transformer and DiffTransformer settings, achieving the best BLEU score of 34.64 and 34.83, respectively. The fine-tuned results show the effectiveness of incorporating ESPFormer into the model architecture.
Table 3: Plug-and-Play and Fine-Tune Boost performance of Transformer and DiffTransformer Baselines on the IWSLT’14 German-to-English dataset, reported as the median over 4 runs. Results with a sign denote the Plug-and-Play performance of a different attention module than the base attention module.
Model Plug-and-Play Fine-Tune Boost ESPFormer 33.38 34.64 Sinkformer 33.36 34.61 Transformer Transformer 33.40 34.61 ESPFormer 33.72 34.83 Sinkformer 33.67 34.81 DiffTransformer DiffTransformer 33.85 34.78

### v2

@@ 2.5 Soft Sorting
Calculating the sliced Wasserstein distance involves evaluating the quantile functions of the distributions, which, in the discrete case, can be boiled down to the sorting operation. Sorting is one of the most common operations in computer science. Yet, the piecewise-linear sorted value function and the integer-valued rank/argsort operators pose a significant obstacle for gradient-based optimization techniques, which are essential in deep learning, as neither of them is differentiable. To incorporate sorting operations into the backpropagation framework, differentiable approximations, known as soft sorting, have been explored. Examples include smoothed rank operators by adding Gaussian noise (Taylor et al., 2008) and by using sigmoid surrogate functions (Qin et al., 2010), parameterizing permutations in terms of a differentiable relaxation (Mena et al., 2018), and relaxing the permutation matrices to be only row-wise stochastic (Grover et al., 2019). Of note, Cuturi et al. (2019) propose a differentiable proxy by viewing sorting as an optimal assignment problem and relaxing it to an entropic optimal transport problem from the input values to an auxiliary probability measure supported on an increasing family of target values.
Prillo & Eisenschlos (2020) introduce a simple yet highly effective continuous relaxation of the argsort operator using the softmax function. Given a vector $v\in\mathbb{R}^{N}$,
@@
approximates the sorting permutation matrix, where the softmax operator is applied row-wise, $d(\cdot,\cdot)$ can be any differentiable semi-metric, and $t$ is a temperature parameter that controls the degree of the approximation.
@@ 3.1 ESP for Uniform Discrete Distributions
Given our application of interest, we specifically focus on uniformly distributed discrete distributions with an equal number of support points, while the method can be readily extended to an unequal number of support points. For a given $N\in\mathbb{N}$, denote the space of uniform discrete distributions with $N$ support as
@@
Consider the 1-dimensional slice of $\mu^{1}$ and $\mu^{2}$ in the $\theta$ direction: $\theta_{\#}\mu^{1}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\theta\cdot x_{i}}$ and $\theta_{\#}\mu^{2}=\frac{1}{N}\sum_{j=1}^{N}\delta_{\theta\cdot y_{j}}$, with $X\theta=[\theta\cdot x_{1},\theta\cdot x_{2},\cdots,\theta\cdot x_{N}]^{T}\in\mathbb{R}^{N}$ and $Y\theta=[\theta\cdot y_{1},\theta\cdot y_{2},\cdots,\theta\cdot y_{N}]^{T}\in\mathbb{R}^{N}$. There exists permutation matrices $A,B\in\mathbf{S}_{N}$ such that $AX\theta\in\mathbb{R}^{N}$ and $BY\theta\in\mathbb{R}^{N}$ are in sorted order, i.e.,
@@
or equivalently, given that $A^{T}=A^{-1}$, $A^{T}B$ represents the transport map from $X\theta$ to $Y\theta$, i.e.,
@@
where $\tau$ denotes the “inverse temperature” hyperparameter, and $\sigma_{l}^{\tau}$ is defined as
@@
$\mathcal{D}^{p}_{p}(\mu^{1},\mu^{2};\theta)=\sum_{i=1}^{N}\sum_{j=1}^{N}\|x_{i}-y_{j}\|^{p}U_{\theta}(\{(x_{i},y_{j})\}),$ (10)
where $U_{\theta}(\{(x_{i},y_{j})\})$ is the $(i,j)^{\prime}$th component of the $U_{\theta}$ matrix. When $\tau=0$, $\sigma_{l}=\frac{1}{L}$ for all $l\in[1,2,\cdots,L]$, then $G$ is simply the mean of all the lifted plans.
@@ 3.2 ESP Attention
To bridge the abstract formulation of the Expected Sliced Plan (ESP) with its practical use in attention mechanisms, we interpret the query and key matrices as empirical probability measures due to their permutation-equivariant nature. Following our notation in Section 2.1, let $W_{Q},W_{K}\in\mathbb{R}^{m\times d},W_{V}\in\mathbb{R}^{d\times d}$ denote the query, key, and value matrices, respectively, and let $X=[x_{1},x_{2},\cdots,x_{N}],x_{i}\in\mathbb{R}^{d},\forall i$ denote the input tokens. Also, let $Q=W_{Q}X\in\mathbb{R}^{m\times N}$ and $K=W_{K}X\in\mathbb{R}^{m\times N}$ denote the queries and keys, respectively. Then, we think of doubly-stochastic attention as a transport plan between the empirical measures $\mu^{Q}$ and $\mu^{K}$, defined as
@@
A key consideration is the choice of slices, $\Theta$. While classic sliced OT uniformly samples slices from $\mathbb{S}^{m-1}$, prior work suggests learning them (Deshpande et al., 2019; Naderializadeh et al., 2021; NaderiAlizadeh et al., 2025). Though this approach is common, learning $\Theta$ introduces additional parameters, increasing the total count in ESP attention and complicating fair comparisons with other mechanisms. Notably, since the input distributions (keys and queries) are themselves learned, optimizing $\Theta$ may be unnecessary. Instead, the distributions can adapt to a fixed slicing scheme, as done in prior work, such as FSPool (Zhang et al., 2020). To avoid extra parameters, in this work, we propose using axis-aligned slices by setting $\Theta=\mathds{I}_{m\times m}$, the identity matrix. This leads to
@@
where $V=W_{V}X\in\mathbb{R}^{d\times N}$ is the value matrix. Figure 1 illustrates the construction of the ESP matrix $G$, which serves as the proposed transport-based attention map, directing how information flows from $V$ to the output. Moreover, Figure 2 compares classic attention (Sinkhorn $S^{0}$), Sinkhorn attention from (Sander et al., 2022) across different iteration counts, and our ESP attention under varying “inverse temperature” hyperparameters. As can be seen, while ESP provides a balanced distribution of attention due to its double stochasticity, the “inverse temperature” parameter controls the number of other tokens each token should pay attention to. The overall pipeline of ESPFormer can be found in Algorithm 1. For completeness, we also report results from explicitly learning the slices and observe no significant benefits compared to axis-aligned slices.
@@
0:  Query matrix $Q\in\mathbb{R}^{m\times N}$, Key matrix $K\in\mathbb{R}^{m\times N}$,Value Matrix $V\in\mathbb{R}^{d\times N}$, SoftSort hyperparameter $t$,and “inverse temperature” hyperparameter $\tau$.
@@ 3.3 Runtime Complexity
ESPFormer begins with query and key projections ($Q=W_{Q}X$ and $K=W_{K}X$), each requiring $\mathcal{O}(mNd)$ operations, where $d$ corresponds to the output dimension. The subsequent SoftSort operation, which involves computing pairwise distances, incurs a complexity of $\mathcal{O}(N^{2})$ per slice. When applied across all $L=m$ slices, this yields a complexity of $\mathcal{O}(mN^{2})$. The computation of transport plans through matrix multiplications with soft permutation matrices contributes an additional $\mathcal{O}(mN^{2})$ operations. The final aggregation of transport plans requires summing $m$ plans of size $N\times N$, also contributing $\mathcal{O}(mN^{2})$ operations. Therefore, the overall runtime complexity of ESPFormer is $\mathcal{O}(mN(N+d))$. In comparison, Sinkhorn’s algorithm exhibits a runtime complexity of $\mathcal{O}((S+m)N^{2})$, where $S$ denotes the number of iterations. A key distinction lies in the parallelization capabilities: while Sinkhorn necessitates sequential processing over $S$ iterations, ESPFormer enables parallel computation across the $m$ slices. This parallelizability allows ESPFormer to scale efficiently with increasing $m$. As demonstrated in Figure 3, ESPFormer offers superior efficiency at all sequence lengths with hard sorting, and surpasses Sinkformer for $S>3$ when using soft sorting. The wall clock runtime analysis of ESPFormer compared to all the baselines can be found in Appendix A.
Figure 3: Runtime analysis of ESPFormer versus Sinkformer with varying iteration counts ($S$) for sequence lengths $N\in\{50,100,500,1000\}$ with $d=1024$, averaged over 10 runs.ESPFormer achieves superior computational efficiency across all sequence lengths with hard sorting, and outperforms Sinkformer for $S>3$ under soft sorting, while maintaining model expressivity. .
@@ 4 Experiments
In this section, we evaluate the performance of ESPFormer across a diverse set of domains, including image classification using ViTs (Dosovitskiy et al., 2021), point cloud classification using set transformers (Lee et al., 2019) and point cloud transformers (Guo et al., 2021), as well as sentiment analysis and neural machine translation via Transformers, highlighting the applicability of our approach in comparison to Differential Transformer (Ye et al., 2025), Sinkformer (Sander et al., 2022), and the Vanilla Transformer (Vaswani et al., 2017). Throughout our experiments, we employed an axis-aligned slicer. our experimental setup and implementation can be found in Appendices B and C.
Table 1: Average and standard deviation (over 3 runs) of ESPFormer’s classification accuracy (%) vs. baselines on the Cats and Dogs dataset under varying data availability. ESPFormer’s performance is reported in three modes: initial soft sort, sharp soft sort, and hard sort.
Data Fraction Baselines ESPFormer
Sinkformer DiffTransformer Transformer Initial Soft Sort Sharp Soft Sort Hard Sort
1% 55.07 $\pm$ 3.34 53.78 $\pm$ 0.28 49.71 $\pm$ 0.31 55.66 $\pm$ 3.95 57.86 $\pm$ 3.77 58.52 $\pm$ 3.73
10% 69.56 $\pm$ 0.32 67.34 $\pm$ 0.11 57.25 $\pm$ 0.22 71.49 $\pm$ 0.43 72.22 $\pm$ 0.37 72.71 $\pm$ 0.36
25% 74.56 $\pm$ 0.58 74.86 $\pm$ 0.17 72.25 $\pm$ 0.16 75.40 $\pm$ 0.38 75.92 $\pm$ 0.31 75.92 $\pm$ 0.28
100% 79.12 $\pm$ 0.17 78.85 $\pm$ 0.11 78.49 $\pm$ 0.09 79.47 $\pm$ 0.12 80.61 $\pm$ 0.11 81.23 $\pm$ 0.11
@@ 4.1 ModelNet40 Classification
The ModelNet40 dataset (Wu et al., 2015) comprises 40 widely recognized 3D object categories and serves as a standard benchmark for point cloud classification. Transformers designed for point clouds and sets have been extensively evaluated on ModelNet40, with notable examples including Set Transformers (Lee et al., 2019) and Point Cloud Transformers (Guo et al., 2021). Table 2 presents the classification results across four runs, comparing different attention mechanisms integrated into Set Transformers and Point Cloud Transformers. Notably, ESPFormer outperforms competing methods, demonstrating superior performance.
Table 2: Test accuracy (%) on the ModelNet40 dataset over 4 runs. Accuracies marked with are reported from (Sander et al., 2022).
@@
Set Transformer 87.8 86.3 85.8 84.7
Set DiffTransformer 89.0 88.7 88.7 88.6
Set Sinkformer 89.1 88.4 88.3 88.1
Set ESPFormer 89.6 89.5 89.4 89.1
Point Cloud Transformer 93.2 92.5 92.5 92.3
Point Cloud DiffTransformer 93.1 92.8 92.7 92.6
Point Cloud Sinkformer 93.1 92.8 92.7 92.5
Point Cloud ESPFormer 93.2 92.9 92.7 92.6
4.2 Sentiment Analysis
We next evaluate ESPFormer on the IMDB dataset (Maas et al., 2011) for sentiment analysis. Following Sander et al. (2022), our architecture comprises an attention‐based encoder followed by a max‐pooling layer trained to classify movie reviews as either positive or negative. Table 3 presents the accuracy improvements achieved by using ESPFormer as the core attention module, compared to baseline models, showcasing its robustness in text classification.
Table 3: Test accuracy (%) for Sentiment Analysis on IMDb.
@@ 4.2 Sentiment Analysis
Transformer 85.30 85.25 85.25 85.20
DiffTransformer 85.50 85.45 85.45 85.40
Sinkformer 85.40 85.39 85.37 85.30
ESPFormer 85.50 85.50 85.47 85.40
We additionally evaluate ESPFormer on the TweetEval sentiment dataset (Barbieri et al., 2020), where tweets are labeled positive or negative. We use the same encoder + max‐pooling architecture and training setup as for IMDb. Table 4 reports the corresponding accuracy metrics.
Table 4: Test accuracy (%) for Sentiment Analysis on TweetEval.
Model Best Median Mean Worst
Transformer 71.50 71.35 71.31 71.10
DiffTransformer 72.60 72.35 72.31 72.00
Sinkformer 72.40 72.30 72.23 71.90
ESPFormer 72.60 72.40 72.36 72.10
@@ 4.3 Neural Machine Translation
We evaluate ESPFormer and Sinkformer using two reference models: the Transformer and its DiffTransformer counterpart (Ye et al., 2025). Both models are trained using the fairseq sequence modeling toolkit (Ott et al., 2019) on the IWSLT’14 German-to-English dataset (Cettolo et al., 2014). The architecture of both models consists of an encoder and a decoder, each with a depth of 6 layers. Initially, we trained both models for 25 epochs using the standard training procedure. After this phase, we performed a Plug-and-Play evaluation, where the attention heads were plugged into the pre-trained models and evaluated on their performance, as shown in Table 5. In this phase, we tested ESPFormer and Sinkformer attention heads, denoted by the respective models in the table, comparing their performance to the base Transformer and DiffTransformer models. Following the Plug-and-Play evaluation, we performed a Fine-Tune Boost phase, where the models were further fine-tuned for an additional 10 epochs. The fine-tuning led to further performance gains, as observed in the table, with ESPFormer showing the highest improvement in both the Transformer and DiffTransformer settings, achieving the best BLEU score of 34.64 and 34.83, respectively. The fine-tuned results show the effectiveness of incorporating ESPFormer into the model architecture.
Table 5: Plug-and-Play and Fine-Tune Boost performance of Transformer and DiffTransformer Baselines on the IWSLT14 German-to-English dataset, reported as the median over 4 runs. Results with a sign denote the Plug-and-Play performance of a different attention module than the base attention module.
Model Plug-and-Play Fine-Tune Boost
Transformer Transformer 33.40 34.61
Sinkformer 33.36 34.61
ESPFormer 33.38 34.64
DiffTransformer DiffTransformer 33.85 34.78
Sinkformer 33.67 34.81
ESPFormer 33.72 34.83
Table 6: Accuracy (%) over three runs (mean $\pm$ standard deviation) across slicer types, slice counts $L$, and inverse temperature $\tau$.
$\boldsymbol{L=1}$ $\boldsymbol{L=8}$ $\boldsymbol{L=32}$ $\boldsymbol{L=64}$ $\boldsymbol{L=128}$
$\boldsymbol{\tau=0.1}$
Learnable $74.30\pm 0.48$ $78.70\pm 0.32$ $79.10\pm 0.22$ $78.40\pm 0.26$ $76.20\pm 0.42$
Frozen $66.50\pm 0.52$ $72.80\pm 0.38$ $78.30\pm 0.18$ $79.20\pm 0.30$ $79.60\pm 0.28$
Axis-Aligned   –   –   – $79.47\pm 0.12$   –
$\boldsymbol{\tau=1.0}$
Learnable $74.30\pm 0.48$ $79.07\pm 0.30$ $78.20\pm 0.35$ $77.80\pm 0.21$ $74.10\pm 0.46$
Frozen $66.50\pm 0.52$ $73.11\pm 0.43$ $77.95\pm 0.25$ $78.80\pm 0.29$ $78.40\pm 0.27$
Axis-Aligned   –   –   – $78.85\pm 0.31$   –
$\boldsymbol{\tau=10}$
Learnable $74.30\pm 0.48$ $79.15\pm 0.29$ $78.06\pm 0.27$ $77.10\pm 0.23$ $74.15\pm 0.44$
Frozen $66.50\pm 0.52$ $73.45\pm 0.41$ $76.85\pm 0.24$ $77.85\pm 0.30$ $78.10\pm 0.26$
Axis-Aligned   –   –   – $77.75\pm 0.32$   –